# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 09 Feb 2024 v2

@@ Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current neural SLAM can achieve ideal results in reconstructing bounded scenes, but this relies on the input of RGB-D images. Neural-based SLAM based only on RGB images is unable to reconstruct the scale of the scene accurately, and it also suffers from scale drift due to errors accumulated during tracking. To overcome these limitations, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to previous work, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has more excellent mapping performance than prior neural SLAM methods, especially in large borderless scenes.
@@ 1 Introduction
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [45] have achieved good results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [41, 42, 37, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately, but it still can not be implemented in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. These problems limit the application and performance of neural-based SLAM systems in large scenes or scenes with numerous objects and texture details. GO-SLAM [42] tries to input RGB images for 3D scene reconstruction, but the result is unsatisfactory. It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which leads to misalignment and performance degradation in large-scale settings. Therefore, these methods are limited to more controlled and smaller room-scale scenes.
Figure 1: Reconstruction results on ScanNet. We present MoD-SLAM, a neural-based monocular dense mapping method. Our method shows a more powerful capability than existing work: GO-SLAM [42].
In this work, we propose MoD-SLAM, a monocular implicit neural-based dense mapping method. In contrast to previous works, our SLAM system no longer relies on RGB-D image inputs. It improves its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [34] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. We also include loop closure detection in the global update so that the system can combat scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
- A novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by using depth estimation
@@ 2 Related work
Monocular Visual SLAM In the realm of robotic navigation and augmented reality, Simultaneous Localization and Mapping (SLAM) plays a pivotal role. Monocular Visual SLAM, relying solely on a single camera, has attracted significant attention due to its cost-effectiveness and applicability in GPS-denied environments. This technique constructs the 3D map of the surroundings while simultaneously estimating the camera’s position and orientation within that map. By processing sequential frames of video, monocular SLAM algorithms detect and track visual features across images, using them to calculate depth and structure from motion. Some works focused on using more efficient data structures like VoxelHashing [5, 27, 20] and Octrees [35, 39] to improve performance. The most notable method in this domain is ORB-SLAM3 [4], a versatile and accurate system that utilizes ORB features to achieve robust tracking and mapping under a wide range of conditions. Another innovative approach is DROID-SLAM [33], which introduces deep learning to enhance tracking accuracy and the overall robustness of the SLAM system. While these methods have set high benchmarks, they are primarily reliant on feature extraction and matching, which can be challenging in feature-poor or dynamic environments.
@@
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking module, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects keyframes that are subsequently fed into the mapping module. 2) The mapping module leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs, which in turn enhances the accuracy of scene reconstruction and concurrently refines the pose precision within the tracking module.
@@ 3 Method
Fig. 2 shows an overview of our method. For an input RGB stream, we employ a two-step process: first, we estimate the depth values of keyframes using a depth estimator, and subsequently, we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). We employ a Gaussian spatial information collection approach within the region to capture both the geometric structure and the visual appearance of the scene. In the case of a scene without distinct boundaries, we employ a contraction function to learn long-range spatial information (Sec. 3.2). Specifically, we achieve loop closure detection to eliminate scene drift (Sec. 3.3). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5).
@@ 3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream will not only be used as a dataset for NeRF sampling (Sec. 3.2), but will also be fed into the pre-trained depth estimation module to generate depth information of the scene and put into a depth distillation module to distill depth information subsequently. The use of priori depth estimation significantly enhances the accuracy of scale in scene reconstructed through monocular mode, ensuring that the dimensions of the scene are precisely rendered.
Depth Estimation Fig. 2 shows the components of our depth estimation module. Our depth estimation module contains two parts: relative depth estimation and metric depth estimation. In the relative depth estimation module, we adopt the training criterion of MiDaS [24] and use the DPT decoder-encoder architecture [23] as backbones for feature extraction to predict the relative depths. In the metric depth estimation module, we follow ZoeDepth [3] by proposing the metric bins module to output the metric depth. The metric bins module consists of five heads, which are combined with the relative depths predicted by DPT through convolutional layers and softmax layers to obtain the final depth prediction.
@@
where $T(k_{m})$ represents the count of $\mathbf{D}_{k_{n}}\in\mathbf{\Theta}_{k_{m}}$, holding $|\mathbf{D}_{k_{m}}-\mathbf{D}_{k_{n}}|<\tau^{\prime}$ true. $\tau^{\prime}$ signifies a minor margin between the current pixel point and its adjacent pixel points. $M$ denotes the aggregate count of samples satisfying the criteria within set $\mathbf{P}$. Our method punishes non-continuous depth values at the surface of the geometric object and avoids bias at the edges of the geometric object.
Training Strategy As mentioned above, to enable our depth estimation module to accurately predict depth, we need to pre-train the MiDaS backbones network to predict relative depth as well as the head of the metric bins module to fine-tune the metric depth.
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Therefore, our model uses multiple metric heads. These different metric heads are used for metric depth prediction in different datasets, which requires us to train different metric heads using different datasets. In addition, because our model incorporates relative depth pretraining, it can also solve the problem of fine-tuning the relative depth backbone in different scenes to some extent.
@@ 3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a fine ray to the center of each pixel point and then performing coarse and fine sampling. This sampling strategy enriches the approach to some extent by integrating all the sampled points on the ray to obtain the color of a pixel. However, it also leads to missing information. This is because the number of sampled points is limited, and the ray fails to encapsulate information within the three-dimensional space surrounding the sampled points. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. To enhance the system’s stability, particularly in scenarios demanding higher accuracy and more sampling points, we address a critical issue. The proximity of neighboring sampling points, denoted as $t_{0}$ and $t_{1}$, often results in a minimal distance $t_{1}-t_{0}$, which can compromise stability during division. Consequently, we introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to bolster the system’s stability. We assume that cone truncations are all symmetric for the cone center axis, then the multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
@@
Figure 3: Scene Reparameterization. To manage unbounded scenes, we use a contract function to map mean and covariance data from spatial sampling (black dashed line) directly to a new range (red solid line). We ignore the data within the sphere of radius 1 (yellow region) while mapping the data from regions beyond a radius of 1 into the spherical coordinate system with a radius of 2 (orange region).
@@
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean function $E(\mathbf{x})$ and covariance function $D(\mathbf{x})$ as:
$\centering f(E(\mathbf{x}),D(\mathbf{x}))=\begin{cases}E(\mathbf{x}):f(\mathbf{\mu})\\ D(\mathbf{x}):\mathbf{J}_{f}(\mathbf{\mu})D(\mathbf{x})\mathbf{J}_{f}(\mathbf{\mu})^{T},\end{cases}\@add@centering$ (8)
our choice for $f$ is relied on Equ. 6.
@@ 3.4 Depth and Color Rendering
Inspired by the recent work in NeRF, we propose to also use Neural radiance filed to reconstruct 3D scenes.
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to world coordinate. Unlike NeRF, since our system is also used to process borderless scenes at the same time, we propose here a more efficient sampling technique, the differential sampling technique. This sampling technique aims to sample more in the near part of the ray and sparsely in the far part. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
To complete the scene reconstruction, we randomly extract $N$ rays and $M$ pixel points in the current frame and keyframes. We input the position set $\mathbf{P}$ to the depth prediction unit in Sec. 3.1 to get the true depth $\mathbf{D}$. We do a multivariate Gaussian variation of the pose set $\mathbf{G}$ in Sec. 3.2 to get the mean as well as the covariance set. The mean and covariance are then passing into the MLPs to obtain the predicted volume density as well as the predicted color. Volume density is used to calculate the occupancy probability $o_{i}$ and the ray termination probability at each sample can then be calculated as $w_{i}=o_{i}\prod_{i=j}^{i-1}(1-o_{j})$. Eventually, the depth and color can be rendered as:
@@ 3.5 Tracking and Mapping
Camera Tracking For more accurate position tracking, in this work, we employ DROID-SLAM [33] as the front end of our system due to its superior position prediction capabilities and minimal position drift. We take the video stream as input, and for each frame maintain two state variables: camera position and inverse depth. The pose set $\left\{\mathbf{G}_{t}\right\}^{N}_{t=0}$ and the inverse depth set $\left\{\mathbf{d}_{t}\right\}^{N}_{t=0}$ are unknown state variables, which are iteratively updated during the inference process of processing a new frame. Each input image is processed through a feature extraction network with a context network. The feature extraction network consists of 6 residual blocks and 3 downsampling layers to generate a dense feature map for each image.
@@ 4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28] and ScanNet [7]. We also conducted a comparative analysis of the reconstruction quality of our system against state-of-the-art SLAM systems and executed ablation studies to validate the soundness of our model’s architecture.
@@ 4.1 Experimental Setup
Ours 0.33 3.21 2.29 29.95 0.862 6
DROID-SLAM[33] 0.42 - 5.03 - - 21
@@
Implementation Details We run our system on a PC with a 4.70 GHz AMD 7900X CPU and NVIDIA RTX 4090 GPU. We use the pre-trained model from DROID-SLAM [33] for tracking and use the same parameters in the tracking process as it. In the reconstruction process, we choose to sample a ray at several points of $N_{strat}=20$ and $N_{imp}=40$. For Replica, we sample $M=300$ pixels, while for Scan-Net, we sample $M=600$ pixels. For loss weights, we choose $\lambda_{c}=1.0$, $\lambda_{dep}=1.0$, $\lambda_{dist}=0.002$, $\lambda_{eik}=0.15$, $\lambda_{sdf}=1.0$. The compared baseline includes classical SLAM systems ORB-SLAM3 [4] and DROID-SLAM [30], newest SLAM systems NICE-SLAM [45], HI-SLAM [41] and GO-SLAM [42] and 3D Gaussian-based SLAM Photo-SLAM [11].

### 21 Feb 2024 v3

@@ Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current neural SLAM can achieve ideal results in reconstructing bounded scenes, but this relies on the input of RGB-D images. NeRF-based SLAM with RGB image input is unable to reconstruct the scale of the scene accurately, and it also suffers from scale drift due to errors accumulated during tracking. To address these issues, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to recent work, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has the state-of-the-art monocular mapping performance, especially in large borderless scenes.
@@ 1 Introduction
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [46] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [42, 43, 38, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still can not be implemented in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. These problems limit the application and performance of neural-based SLAM systems in large scenes or scenes with numerous objects and texture details. GO-SLAM [43] inputs RGB images for 3D scene reconstruction, but the result is unsatisfactory. It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which results in accumulated positional drift, leading to inaccuracies in tracking and mapping process.
Figure 1: Reconstruction results on ScanNet 0207 [7]. We present MoD-SLAM, a neural-based monocular dense mapping method. Our method shows a more powerful capability than existing work: GO-SLAM [43].
In this work, we propose MoD-SLAM, a monocular implicit neural-based dense mapping method. In contrast to previous works, our SLAM system no longer relies on RGB-D image inputs. We improve its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [35] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. We also include loop closure detection in the global update so that the system can combat scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
- A novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by using depth estimation.
@@ 2 Related work
Monocular Visual SLAM In the realm of robotic navigation and augmented reality, Simultaneous Localization and Mapping (SLAM) plays a pivotal role. Monocular Visual SLAM has attracted significant attention due to its cost-effectiveness compared to stereo SLAM. This technique constructs the 3D map of the surroundings while simultaneously estimating the camera’s position and orientation within that map. By processing sequential frames of video, monocular SLAM algorithms detect and track visual features across images, using them to calculate depth and structure from motion. Some works focused on using more efficient data structures like VoxelHashing [5, 27, 20] and Octrees [36, 40] to improve performance. The most notable method in this domain is ORB-SLAM3 [4], a versatile and accurate system that utilizes ORB features to achieve robust tracking and mapping under a wide range of conditions. Another innovative approach is DROID-SLAM [34], which introduces deep learning to enhance tracking accuracy and the overall robustness of the SLAM system. While these methods have set high benchmarks, they are primarily reliant on feature extraction and matching, which can be challenging in feature-poor or dynamic environments.
@@
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking module, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects keyframes that are subsequently fed into the mapping module. 2) The mapping module leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs.
@@ 3 Method
For an input RGB stream, we employ a two-step process: first, we estimate the depth values of keyframes using a depth estimator, and subsequently, we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). We employ a Gaussian spatial information collection approach within the region to capture both the geometric structure and the visual appearance of the scene. In the case of a scene without distinct boundaries, we employ a contraction function to deal with long-range spatial information (Sec. 3.2). Specifically, we achieve loop closure detection to eliminate scene drift (Sec. 3.3). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5). Fig. 2 shows an overview of our method.
@@ 3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth data. The use of priori depth estimation significantly enhances the accuracy of scale in scene reconstructed through monocular mode. Our depth estimation unit is inspired by ZoeDepth [3].
Depth Estimation Fig. 2 shows the components of our depth estimation module. Our depth estimation module contains two parts: relative depth estimation and metric depth estimation. In the relative depth estimation module, we adopt the training criterion of MiDaS [24] and use the DPT decoder-encoder architecture [23] as backbones for feature extraction to predict the relative depths. In the metric depth estimation module, we follow ZoeDepth [3] by proposing the metric bins module to output the metric depth.
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. There is also not a linear multiplicative relationship between the scale predicted by ZoeDepth and the scale predicted by the mapping process. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth information predicted by our depth estimation module.
@@
where $T(k_{m})$ represents the count of $\mathbf{D}_{k_{n}}\in\mathbf{\Theta}_{k_{m}}$, holding $|\mathbf{D}_{k_{m}}-\mathbf{D}_{k_{n}}|<\tau^{\prime}$ true. $\tau^{\prime}$ signifies a minor margin between the current pixel point and its adjacent pixel points. $M$ denotes the aggregate count of samples satisfying the criteria within set $\mathbf{P}$.
@@ 3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a fine ray to the center of each pixel point and then performing coarse and fine sampling. This approach leads to missing information. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. The proximity of neighboring sampling points, denoted as $t_{0}$ and $t_{1}$, often results in a minimal distance $t_{1}-t_{0}$, which can compromise stability during division. Consequently, we introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
@@
Figure 3: Scene Reparameterization. To manage unbounded scenes, we use a contract function to map mean and covariance data from spatial sampling (black dashed line) directly to a new range (red solid line). We save the data within the sphere of radius 1 (yellow region) while mapping the data from regions beyond a radius of 1 into the spherical coordinate system with a radius of 2 (orange region).
@@
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean function $E(\mathbf{x})$ and covariance function $D(\mathbf{x})$. our choice for $f$ is relied on Equ. 6.
$\centering f(E(\mathbf{x}),D(\mathbf{x}))=\begin{cases}E(\mathbf{x}):f(\mathbf{\mu})\\ D(\mathbf{x}):\mathbf{J}_{f}(\mathbf{\mu})D(\mathbf{x})\mathbf{J}_{f}(\mathbf{\mu})^{T}.\end{cases}\@add@centering$ (8)
@@ 3.4 Depth and Color Rendering
Inspired by the recent work in NeRF, we propose to also use Neural radiance field to reconstruct 3D scenes.
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, since our system is also used to process borderless scenes at the same time, we propose here a more efficient sampling technique, the differential sampling technique. This sampling technique aims to sample more in the near part of the ray and sparsely in the far part. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
To complete the scene reconstruction, we randomly extract $N$ rays and $M$ pixel points in the current frame and keyframes. We input the position set $\mathbf{P}$ to the depth prediction module in Sec. 3.1 to get the true depth $\mathbf{D}$. We do a multivariate Gaussian variation of the pose set $\mathbf{G}$ in Sec. 3.2 to get the mean as well as the covariance set. The mean and covariance are then passing into the MLPs to obtain the predicted volume density as well as the predicted color. Volume density is used to calculate the occupancy probability $o_{i}$ and the ray termination probability at each sample can then be calculated as $w_{i}=o_{i}\prod_{i=j}^{i-1}(1-o_{j})$. Eventually, the depth and color can be rendered as:
@@ 3.5 Tracking and Mapping
Camera Tracking For more accurate position tracking, in this work, we employ DROID-SLAM [34] as the front end of our system due to its superior position prediction capabilities and minimal position drift. We take the RGB stream as input, and for each frame maintain two state variables: camera position and inverse depth. The pose set $\left\{\mathbf{G}_{t}\right\}^{N}_{t=0}$ and the inverse depth set $\left\{\mathbf{d}_{t}\right\}^{N}_{t=0}$ are unknown state variables, which are iteratively updated during the inference process of processing a new frame. Each input image is processed through a feature extraction network with a context network.
@@
Figure 5: Reconstruction results on ScanNet [7]. We show the full view of the reconstructed 3D scenes. Our system can better reconstruct the scale of the scenes.
@@ 4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28] and ScanNet [7]. We also conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems.
@@ 4.1 Experimental Setup
Implementation Details We run our system on a PC with a 4.70 GHz AMD 7900X CPU and NVIDIA RTX 4090 GPU. We use the pre-trained model from DROID-SLAM [34] for tracking and use the same parameters in the tracking process as it. In the reconstruction process, we choose to sample a ray at several points of $N_{strat}=20$ and $N_{imp}=40$. For loss weights, we choose $\lambda_{c}=1.0$, $\lambda_{dep}=1.0$, $\lambda_{dist}=0.002$, $\lambda_{eik}=0.15$, $\lambda_{sdf}=1.0$. The compared baseline includes classical SLAM systems ORB-SLAM3 [4] and DROID-SLAM [30], newest SLAM systems NICE-SLAM [46], HI-SLAM [42] and GO-SLAM [43] and 3D Gaussian-based SLAM Photo-SLAM [11].
@@
Ours 0.33 3.11 2.13 30.95 0.882 6
DROID-SLAM[34] 0.42 17.59 5.03 7.68 0.191 21